The mass of Sputnik I was 83.5 kg, and its distances from the center of the Earth at apogee and perigee were 7330 km and 6610 km, respectively.
Find the difference in gravitational potential energy for Sputnik I as it moved from apogee to perigee.
Answer #1

G = 6.67 x 10-11 N-m²/kg²
ME = 5.97 x 1024 kg
U = -G ME m / r
ΔU = Up - Ua = (-G ME m / rp) - (-G ME m / ra) = G ME m (1 / ra - 1 / rp)

 

     = (6.67*10-11)*(5.97*1024)*83.5((1/7330*103)-(1/6610*103))

 

     = - 4.94 x 108 J